We got demand function q=6000-30p (q quantity, p=price). The total cost function is C(q)=72000+60q a) Solve the demand function for p and define its domain b) Find the marginal cost funtion MC and explain the result c) Find the income function R(q)
1
Expert's answer
2020-05-25T17:56:47-0400

a) "p=200-1\/30q"

b) "MC = dC\/dq= 60."

c)"R= P*Q-C(q)= q(200-1\/30q) - 72000-60q=\n140q-1\/30q^2-72000"